OpenStudy (phebe):

MEDALSSSSS!!!!! x^2-5x-36 factor this trinoial

OpenStudy (anonymous):

x^2 +5*x -36 = 0 Separate : ( x^2 -4*x ) + ( 9*x -36 ) = 0 Commutative Law : ( x^2 +9*x ) + ( -4*x -36 ) = 0 Distributive Law : x*( x +9 ) + -4*( x +9 ) = 0 Factor : ( x -4 )*( x +9 ) Answer : C.(x+9)(x-4)